OpenStudy (anonymous):

the denstiy of mercury at 0.00c is 13600 kgm^-3. what is the density of mercury at 50c

OpenStudy (anonymous):

If you knew the coefficient of thermal expansion (linear) you could cube it and divide the density at 0o by that factor, as expansion will lower density.

OpenStudy (anonymous):

not linear.. Volume expansion co efficient :P liquids don't have linear expansion co efficient.. and its not cubing .. its almost 3 times. and only works for solids!

OpenStudy (anonymous):

No, @Mashy, you find the bulk expansion coefficient by cubing (1 + dx) where dx comes from the linear expansion coefficient. For dx<<1, (1+dx)^n = approx. (1+n dx) +... higher-order terms much smaller. I should have been clearer in my explanation, however, but we like to fall short of just giving the answer.

OpenStudy (anonymous):

The expansion works for any volume, liquid or solid. For gases, we go another route.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Thermal_expansion see the table mate.. the volume expansion coefficient is obtained by tripling the linear expansion coefficient

OpenStudy (anonymous):

Yes, but it is only an approximation, as I have noted. We are both correct enough.

OpenStudy (anonymous):

yea.. the approximation itself is 3 times .. not CUBE :P..

OpenStudy (anonymous):

(L+ dx)^3 = actual volume change (L+ dx)^3 = L(1 + 3 dx/L + higher order terms in the expansion) (L + dx) = approximately L(1 + 3 dx/L) which is where the 3 comes from in the bulk expansion coefficient, unless the volume is what was measured when they determined the coefficient, in which case no 3, just dV/V..